7. Property taxes for an Ontario resident over the years are listed.

Year: 2007 2008 2009 2010 2011 2012

Taxes: $2,900 $3,000 $3,150 $3,300 $3,400 $3,600

b) Calculate the least squares equation. (4 marks)
c) Predict the property tax for the year 2020. (2 marks)
1
Expert's answer
2020-04-22T19:32:26-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 X & Y \\ \hline
 0 & 2900\\
 1 & 3000 \\
 2 & 3150 \\
 3 & 3300 \\
 4 & 3400 \\
 5 & 3600
\end{array}"

"\bar{x}={\textstyle\sum_{i=1}^nx_i \over n} ={0+1+2+3+4+5\over 6}=2.5"

"\bar{y}={\textstyle\sum_{i=1}^ny_i \over n} ={2900+3000+3150+3300+3400+3600\over 6}=3225"

"\textstyle\sum_{i=1}^nx_i =15"

"\textstyle\sum_{i=1}^nx_i ^2 =55"


"\textstyle\sum_{i=1}^ny_i =19350"


"\textstyle\sum_{i=1}^ny_i ^2=62742500"

"\textstyle\sum_{i=1}^nx_iy_i =50800"


"S_{xx}=\textstyle\sum_{i=1}^n(x_i -\bar{x})^2=\textstyle\sum_{i=1}^nx_i ^2-n\bar{x}^2=17.5"

"S_{yy}=\textstyle\sum_{i=1}^n(y_i -\bar{y})^2=\textstyle\sum_{i=1}^ny_i ^2-n\bar{y}^2=338750"

"S_{xy}=\textstyle\sum_{i=1}^n(x_i-\bar{x})(y_i -\bar{y})="

"=\textstyle\sum_{i=1}^nx_iy_i -n\bar{x}\bar{y}=2425"

b. Trend line: "y=A+Bx"


"B={S_{xy} \over S_{xx}}={2425 \over 17.5}\approx138.57142857"


"A=\bar{y}-B\bar{x}=3225-{2425 \over 17.5}\cdot2.5\approx2878.57142857"

"y=2878.5714+138.5714x"

"x" is the number of years from 2007.


c. "2020-2007=13"


"y=2878.5714+138.5714(13)\approx4680"

The property tax for the year 2020 will be $4680.